The mean annual cost to attend a college is $44,324. Assume the distribution of  annual costs follows the normal probability distribution and the standard deviation is $4,590.

Ninety percent of all students pay less than what amount?


1
Expert's answer
2021-02-03T03:44:34-0500

Let "X=" the annual cost to attend a college: "X\sim N(\mu, \sigma^2)"

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=\$44324, \sigma=\$4590"


"P(X<x)=P(Z<\dfrac{X-44324}{4590})=0.9"

"\dfrac{X-44324}{4590}=1.28155157"

"X=50206.32"

Ninety percent of all students pay less than $50206.32